# Privacy-aware Berrut Approximated Coded Computing for Federated Learning

### v1

@@ Abstract
Federated Learning (FL) is an interesting strategy that enables the collaborative training of an AI model among different data owners without revealing their private datasets. Even so, FL has some privacy vulnerabilities that have been tried to be overcome by applying some techniques like Differential Privacy (DP), Homomorphic Encryption (HE), or Secure Multi-Party Computation (SMPC). However, these techniques have some important drawbacks that might narrow their range of application: problems to work with non-linear functions and to operate large matrix multiplications and high communication and computational costs to manage semi-honest nodes. In this context, we propose a solution to guarantee privacy in FL schemes that simultaneously solves the previously mentioned problems. Our proposal is based on the Berrut Approximated Coded Computing, a technique from the Coded Distributed Computing paradigm, adapted to a Secret Sharing configuration, to provide input privacy to FL in a scalable way. It can be applied for computing non-linear functions and treats the special case of distributed matrix multiplication, a key primitive at the core of many automated learning tasks. Because of these characteristics, it could be applied in a wide range of FL scenarios, since it is independent of the machine learning models or aggregation algorithms used in the FL scheme. We provide analysis of the achieve privacy and complexity of our solution and, due to the extensive numerical results performed, it can be observed a good trade-off between privacy and precision.
@@ I Introduction
In this paper, we exclusively focus on the privacy concerns that have been highlighted in the literature [2, 3], such as the possibility of membership inference attacks [4, 5], inferring class representatives [6], inferring properties [7] or even being able to reconstruct the original data by inverting gradients [8]. These concerns have led to the emergence of different privacy enhancing technologies (PET) focused on trying to avoid privacy leaks [9], being the most used strategies the following ones:
@@
Within this context, we propose to face the problem from the perspective of a large-scale distributed computing system, since FL naturally matches this scheme: a complex computation (global learning model) is distributedly performed by several worker nodes. Anyway, large-scale distributed computing systems have to deal with two important issues: communication overheads and straggling nodes. The former is due to the exchange of intermediate results in order to be able to collaboratively obtain the final one. The latter is due to slow worker nodes that reduce the runtime of the whole computation. Preciselly, the Coded Distributed Computing (CDC) paradigm [22, 23] arose to reduce both problems by combining coding techniques and redundant computation approaches. However, the CDC techniques cannot simultaneously guarantee privacy and work with non-linear functions. Besides, CDC schemes are usually applied for decentralized computations (one master node and one input dataset), but for FL (multiple datasets) it is more appropriate to work with a scheme like secret sharing protocols [24, 25], whose most famous approach is the Shamir Secret Sharing (SSS) [26]. In this context, we propose a CDC-based solution with the following contributions:
@@
The rest of the paper is organized as follows. Section II briefly reviews some relevant works related to our proposal, and includes a description of the BACC approach [27]. Section III describes our first contribution PBACC and includes an analysis of the achieved privacy. In Section IV, we detail how to extend PBACC to a multi-input secret sharing setting (PBSS) and the analysis its computational complexity. In Section V we detail our contribution to extend our scheme for matrix multiplication, including the privacy and complexity analysis. In Section VI we test our proposal and, finally, Section VII summarizes the main conclusions of this work, and outlines some future work.
@@ II Related Work
All the previously mentioned CDC schemes are suitable for computations with one master node and one input data set. In FL settings, where there are multiple input data sets, one per node (worker), it would be more adequate to use the philosophy of the polynomial sharing approach [34], based on the Shamir Secret sharing (SSS) [26] approach. In SSS the private information (secret) is split up in parts or shares, which are sent to the workers. The secret cannot be reassembled unless a sufficient number of nodes work together to reconstruct the original information.
@@ III-A Preliminaries
Given an arbitrary matrix function $f:\mathbb{V}\rightarrow\mathbb{U}$ over some input data $\mathbf{X}=(X_{0},\dots,X_{K-1})$, where $\mathbb{U}$ and $\mathbb{V}$ are vector spaces of real matrices, BACC performs the approximate evaluation of $f(\mathbf{X})=\bigl{(}f(X_{1}),\dots,f(X_{K-1})\bigr{)}$, where $X_{i}\in\mathbb{V}$, in a numerically stable form and with a bounded error. The computation is executed in a decentralized configuration with one master node, who owns the data, and $N$ worker nodes in charge of computing $f(\cdot)$ at some interpolation points which encode $\mathbf{X}$ in a form that prevents the workers to learn much information about the data. The general procedure is depicted in Fig. 1, and explained next.
@@ III-A4 Decoding and recovery
where $\tilde{z}_{i}\in\mathcal{S}$ are the interpolation points $\mathcal{S}=\{\cos\frac{j\pi}{N-1},j\in\mathcal{F}\}$ from the $n$ faster nodes. At this moment, the master node computes the approximation $f(X_{i})\approx r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{i})$, for $i\in\{0,\dots,K-1\}$.
@@ III-B PBACC: Adding privacy to BACC
$w(z)$ $=\sum_{i=0}^{K-1}\frac{\frac{(-1)^{j}}{(z-\alpha_{i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}X_{i}+\sum_{i=0}^{T-1}\frac{\frac{(-1)^{K+i}}{(z-\alpha_{K+i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}R_{i}$ (2)
$=\sum_{i=0}^{K+T-1}\frac{\frac{(-1)^{j}}{(z-\alpha_{i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}W_{i},$
where $\mathbf{W}=(X_{0},\dots,X_{K-1},R_{0},\dots,R_{T-1})=(W_{0},\dots,W_{K+T-1})$. Here, $\{R_{0}:i=1,\dots,T-1\}$ are random data points independently generated according to a given distribution $\mathcal{G}$ to be specified later. To be able to accomplish the interpolation of the values of $f(X_{i})$ for all $i$, we propose to select as interpolation points, the Chebyshev points of first kind, and as evaluation points the shifted points $\beta_{j}=b+\alpha_{j}=b+\operatorname{Re}(e^{\imath(2j+1)\pi/(2T)})$ where $b\in\mathbb{R}$, for $j=0,\dots,T-1$. Clearly, adding the randomness coefficients in (2) does not modify the identities $w(\alpha_{i})=X_{i}$ for $0\leq i<K$. The proposed scheme will be referred to as Privacy-aware Berrut Approximated Coded Computing (PBACC) hereafter.
@@ III-C Measuring the privacy of PBACC
Our focus is this paper is not on perfect information-theoretical privacy, where the adversaries cannot learn anything about the local data sets of honest clients, but on bounded information leakage. Therefore, similar to other works [33, 40], we use the mutual information $I(\mathbf{X};\mathbf{Y})$, which is a measure of the statistical correlation of two random variables $\mathbf{X}$ and $\mathbf{Y}$, as the privacy leakage metric. One of the variables $\mathbf{X}$ will correspond to the private input, while the other $\mathbf{Y}$ will represent the encoded output, and is a worst-case for the information accessible to the curious or semi-honest colluding nodes. Specifically, we assume that, in our system, up to $T$ nodes can be honest but curious, or semi-honest in short. This means that these nodes, whose identities are unknown to the remaining honest nodes, can attempt to disclose information on the private data $\mathbf{X}$ by means of observation $\mathbf{Y}$ of the encoded information and direct message exchange among them. Notice that semi-honest nodes respect the computation protocol, i.e., they do not inject malicious or malformed messages to deceive the master node.
Generally, the exact computation of $I(\mathbf{X};\mathbf{Y})$ under the described threat model is not feasible. Instead, since our goal is to obtain a bounded leakage assurance, we leverage the existent results about the capacity of a Multi-Input Multi-Output (MIMO) channel under some specific constraints [41], to bound the mutual information, using similar techniques to [33]. Such bound is based in the fact that a MIMO channel with $K$ transmitter antennas and $T$ receiver antennas is, conceptually from an information-theoretic perspective, equivalent to a signal composed of the $K$ encoded elements of the input in BACC, and an observation cooperatively formed by the $T$ colluding nodes, which collect and possibly process these encoded messages to infer information about the data. In this configuration, the encoded noise of the channel acts as the element that provides security, since it reduces the amount of information received by the colluding nodes. We assume for the analysis that follows that the noise is Gaussian (so the channel is an AWGN vector channel).
@@
is one of the rational interpolating terms. Without loss of generality, we assume that each entry of $X_{i}$ is a realization of a random variable with a distribution supported in the interval $D_{s}\triangleq[-s,s]$, for all $i=\{0,\dots,K-1\}$, and that the random coefficients $R_{i}\sim\mathcal{N}(0,\frac{\sigma_{n}^{2}}{T})$, where $\mathcal{N}(0,a)$ denotes the Gaussian distribution with zero mean and variance $a$. Note that, having a finite support $D_{s}$, $X_{i}$ has finite power. If we denote $\mathbf{Y}=(w(z_{i_{1}}),w(z_{i_{2}}),\dots,w(z_{i_{T}}))^{T}$, for the messages observed by the $T$ colluding adversaries $\mathcal{T}=\{i_{1},i_{2},\dots,i_{T}\}$, where $z_{i_{h}}$ are some of the Chebyshev points of second kind, we can write
$\mathbf{Y}=Q_{T}\mathbf{X}+\tilde{Q}_{T}\mathbf{R},$ (3)
where the matrices $Q_{T}$ and $\tilde{Q}_{T}$ are given by
$Q_{T}\triangleq\begin{pmatrix}q_{0}(z_{i_{1}})&\ldots&q_{K-1}(z_{i_{1}})\\ q_{0}(z_{i_{2}})&\ldots&q_{K-1}(z_{i_{2}})\\ \vdots&\ddots&\vdots\\ q_{0}(z_{i_{T}})&\ldots&q_{K-1}(z_{i_{T}})\end{pmatrix}_{T\times K},$
$\tilde{Q}_{T}\triangleq\begin{pmatrix}q_{K}(z_{i_{1}})&\ldots&q_{K+T-1}(z_{i_{1}})\\ q_{K}(z_{i_{2}})&\ldots&q_{K+T-1}(z_{i_{2}})\\ \vdots&\ddots&\vdots\\ q_{K}(z_{i_{T}})&\ldots&q_{K+T-1}(z_{i_{T}})\end{pmatrix}_{T\times T}.$
The amount of information revealed to $\mathcal{T}$ —the information leakage $I_{L}$— is defined in this work as the worst-case achievable mutual information for the colluding nodes [40]
$I_{L}\triangleq\max_{\mathcal{T}}\sup_{P_{\mathbf{X}}:|X_{i}|\leq s|,\forall i\in[K]}I(\mathbf{Y};\mathbf{X}),$
where $P_{\mathbf{X}}$ is the probability density function of $\mathbf{X}$, and the maximization applies to any $\mathcal{T}\subset[N]$. As $\|X_{i}\|\leq s$, this implies that the power $\mathbb{E}[\|X_{i}\|]^{2}\leq s^{2}$. Therefore, the latter equation can be re-written as
$I_{L}\leq\max_{\mathcal{T}}\sup_{P_{\mathbf{X}}:\mathbb{E}[\|X_{i}\|^{2}]\leq s^{2}}I(\mathbf{Y};\mathbf{X}).$ (4)
In order to bound $I_{L}$, since the noise used for privacy in BACC is Gaussian, we consider an equivalent formulation of (3) as a MIMO channel with $K$ transmitter antennas and $T$ receiver antennas, so the input-output relation can be defined as
@@
where $H_{T\times K}$ is the channel gain matrix known by both transmitter and receiver, and $\mathbf{Z}_{T\times 1}$ is a vector of additive Gaussian noise with 0 mean. Further, let us denote the covariance matrix of the vector $\mathbf{Z}$ as $\Sigma_{\mathbf{Z}}$. Using known results on the capacity of a MIMO channel with the same power allocation constraint and correlated noise [41],
$C=\sup_{P_{\mathbf{X}}}I(\mathbf{Y};\mathbf{X})=\log_{2}|I_{T}+PH\Sigma^{-1}_{\mathbf{Z}}H^{\dagger}|,$
where $P$ is the maximum power of each transmitter antenna, $I_{T}$ is the identity matrix of order $T$ and ||$\cdot$|| ⋅ | is the determinant of a matrix. Thus, using (4) and assuming that the noise is uncorrelated, we have
$\ I_{L}\leq\max_{\mathcal{T}}\log_{2}\bigl{|}I_{T}+\frac{s^{2}T}{\sigma^{2}_{n}}\tilde{\Sigma}_{T}^{-1}\Sigma_{T}\bigr{|},$
where $\tilde{\Sigma}_{T}\triangleq\tilde{Q}_{T}\tilde{Q}_{T}^{\dagger}$ and $\Sigma_{T}\triangleq Q_{T}Q_{T}^{\dagger}$.
Since the information leakage grows with $K$, the privacy metric $\imath_{L}$ of BACC is defined in this paper as the normalized value of $I_{L}$, namely $\imath_{L}=\frac{I_{L}}{K}$, and we will say that BACC has $\epsilon$-privacy if $\imath_{L}<\epsilon$, where $\epsilon>0$ is the target security parameter of the scheme and represents the maximum amount of leaked information per data point that is allowed for a fixed number of colluding nodes $T$. The value of this parameter will strongly depend on the specific ML model or application under consideration at each case.

### v2

@@ Abstract
Federated Learning (FL) is an interesting strategy that enables the collaborative training of an AI model among different data owners without revealing their private datasets. Even so, FL has some privacy vulnerabilities that have been tried to be overcome by applying some techniques like Differential Privacy (DP), Homomorphic Encryption (HE), or Secure Multi-Party Computation (SMPC). However, these techniques have some important drawbacks that might narrow their range of application: problems to work with non-linear functions and to operate large matrix multiplications and high communication and computational costs to manage semi-honest nodes. In this context, we propose a solution to guarantee privacy in FL schemes that simultaneously solves the previously mentioned problems. Our proposal is based on the Berrut Approximated Coded Computing, a technique from the Coded Distributed Computing paradigm, adapted to a Secret Sharing configuration, to provide input privacy to FL in a scalable way. It can be applied for computing non-linear functions and treats the special case of distributed matrix multiplication, a key primitive at the core of many automated learning tasks. Because of these characteristics, it could be applied in a wide range of FL scenarios, since it is independent of the machine learning models or aggregation algorithms used in the FL scheme. We provide analysis of the achieved privacy and complexity of our solution and, due to the extensive numerical results performed, a good trade-off between privacy and precision can be observed.
@@ I Introduction
In this paper, we exclusively focus on the privacy concerns that have been highlighted in the literature [2, 3], such as the possibility of membership inference attacks [4, 5], inferring class representatives [6], inferring properties [7] or even being able to reconstruct the original data by inverting gradients [8]. These concerns have led to the emergence of different privacy enhancing technologies (PETs) focused on trying to avoid privacy leaks [9], being the most used strategies the following ones:
@@
Within this context, we propose to face the problem from the perspective of a large-scale distributed computing system, since FL naturally matches this scheme: a complex computation (global learning model) is distributedly performed by several worker nodes. Anyway, large-scale distributed computing systems have to deal with two important issues: communication overheads and straggling nodes. The former is due to the exchange of intermediate results in order to be able to collaboratively obtain the final one. The latter is due to slow worker nodes that reduce the runtime of the whole computation. Precisely, the Coded Distributed Computing (CDC) paradigm [22, 23] arose to reduce both problems by combining coding techniques and redundant computation approaches. However, the CDC techniques cannot simultaneously guarantee privacy and work with non-linear functions. Besides, CDC schemes are usually applied for decentralized computations (one master node and one input dataset), but for FL (multiple datasets) it is more appropriate to work with a scheme like secret sharing protocols [24, 25], whose most famous approach is the Shamir Secret Sharing (SSS) [26]. In this context, we propose a CDC-based solution with the following contributions:
@@
The rest of the paper is organized as follows. Section II briefly reviews some relevant works related to our proposal, and includes a description of the BACC approach [27]. Section III describes our first contribution PBACC and includes an analysis of the achieved privacy. In Section IV, we detail how to extend PBACC to a multi-input secret sharing setting (PBSS), the analysis of its computational complexity, and a comparison with other SMPC and HE based approaches. In Section V, we detail our contribution to extend our scheme for matrix multiplication, including the privacy and complexity analysis, and a comparison with other secure matrix multiplication approaches. In Section VI, we test our proposal and, finally, Section VII summarizes the main conclusions of this work, and outlines some future work.
@@ II Related Work
All the previously mentioned CDC schemes are suitable for computations with one master node and one input data set. In FL settings, where there are multiple input data sets, one per node (worker), it would be more adequate to use the philosophy of the polynomial sharing approach [34], based on the Shamir Secret sharing (SSS) [26] approach. In SSS, the private information (secret) is split up in parts or shares, which are sent to the workers. The secret cannot be reassembled unless a sufficient number of nodes work together to reconstruct the original information.
@@ III-A Preliminaries
Given an arbitrary matrix function $f:\mathbb{V}\rightarrow\mathbb{U}$ over some input data $\mathbf{X}=(X_{0},\dots,X_{K-1})$, where $\mathbb{U}$ and $\mathbb{V}$ are vector spaces of real matrices, BACC performs the approximate evaluation of $f(\mathbf{X})=\bigl{(}f(X_{1}),\dots,f(X_{K-1})\bigr{)}$, where $X_{i}\in\mathbb{V}$, in a numerically stable form and with a bounded error. The computation is executed in a decentralized configuration with one master node, who owns the data, and $N$ worker nodes in charge of computing $f(\cdot)$ at some interpolation points which encode $\mathbf{X}$. The general procedure is depicted in Fig. 1, and explained next.
@@ III-A4 Decoding and recovery
where $\tilde{z}_{i}\in\mathcal{S}$ are the evaluation points $\mathcal{S}=\{\cos\frac{j\pi}{N-1},j\in\mathcal{F}\}$ from the $n$ faster nodes. At this moment, the master node computes the approximation $f(X_{i})\approx r_{\mathrm{Berrut},\mathcal{F}}(\alpha_{i})$, for $i\in\{0,\dots,K-1\}$.
Threat Model. We assume that, in our system, from the $N$ total worker nodes, up to $c$ nodes can be honest but curious, or semi-honest in short. This means that these nodes, whose identities are unknown to the remaining honest nodes, can attempt to disclose information on the private data, denoted by $\mathbf{X}$, by means of observation of the encoded information received $\mathbf{Y}$ and direct message exchange among them. Notice that semi-honest nodes respect the computation protocol, i.e., they do not inject malicious or malformed messages to deceive the master node.
Since we are dealing with rational functions, our scheme cannot achieve perfect information-theoretical privacy, where the adversaries cannot learn anything about the local data sets of honest clients. Instead, the focus is to achieve a bounded information leakage lower than $\epsilon$, the target security parameter of the scheme, which represents the maximum amount of leaked information per data point that is allowed for a fixed number of colluding nodes $c$. The value of this parameter will strongly depend on the specific ML model or application under consideration at each use case.
@@ III-B PBACC: Adding privacy to BACC
$w(z)$ $=\sum_{i=0}^{K-1}\frac{\frac{(-1)^{i}}{(z-\alpha_{i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}X_{i}+\sum_{i=0}^{T-1}\frac{\frac{(-1)^{K+i}}{(z-\alpha_{K+i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}R_{i}=\sum_{i=0}^{K+T-1}\frac{\frac{(-1)^{i}}{(z-\alpha_{i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}}W_{i},$ (2)
where $\mathbf{W}=(X_{0},\dots,X_{K-1},R_{0},\dots,R_{T-1})=(W_{0},\dots,W_{K+T-1})$. Here, $\{R_{i}:i=1,\dots,T-1\}$ are random data points independently generated according to a given distribution $\mathcal{G}$ to be specified later. To be able to accomplish the interpolation of the values of $f(X_{i})$ for all $i$, we propose to select as interpolation points of $X$, the Chebyshev points of first kind, and as interpolation points of $R$, the shifted Chebyshev points of first kind $\beta_{j}=b+\alpha_{j}=b+\operatorname{Re}(e^{\imath(2j+1)\pi/(2T)})$ where $b\in\mathbb{R}$, for $j=0,\dots,T-1$.
Re-writing eq. (2) as
$w(z)=\sum_{i=0}^{K-1}X_{i}q_{i}(z)+\sum_{i=K}^{K+T-1}R_{i-K}q_{i}(z),$
where $q_{i}(z)$ is:
$q_{i}(z)=\frac{\frac{(-1)^{i}}{(z-\alpha_{i})}}{\sum_{j=0}^{K+T-1}\frac{(-1)^{j}}{(z-\alpha_{j})}},$
we can easily check that $q_{i}(z)=1$ when $z=\alpha_{i}$, and $q_{i}(z)=0$ when $z=\alpha_{j}$ for all $j\neq i$. This demonstrates the identity $w(\alpha_{i})=X_{i}$ for $i=\{0,1,\dots,K-1\}$, and $w(\alpha_{j})=R_{j}$ for $j=\{K,K+1,\dots,K+T-1\}$. Thus, the Berrut rational interpolation can be used for interpolating the values of $X_{i}$.
The proposed scheme will be referred to as Privacy-aware Berrut Approximated Coded Computing (PBACC) hereafter.
In the way we have added privacy, PBACC inherits the features of the original BACC (dealing with non-linear functions and resisting stragglers) at a very low cost on precission loss. Our experimentation on the scheme (Section VI) demonstrates this behaviour.
@@ III-C Measuring the privacy of PBACC
As stated in the threat model, our focus is not on perfect information-theoretical privacy but on bounded information leakage. Therefore, similar to other works [33, 40], we use the mutual information $I(\mathbf{X};\mathbf{Y})$, which is a measure of the statistical correlation of two random variables $\mathbf{X}$ and $\mathbf{Y}$, as the privacy leakage metric. One of the variables $\mathbf{X}$ will correspond to the private input, while the other $\mathbf{Y}$ will represent the encoded output, and is a worst-case for the information accessible to the curious or semi-honest colluding nodes. Specifically, we assume that, in our system, up to $c$ nodes can be honest but curious, or semi-honest in short.
Generally, the exact computation of $I(\mathbf{X};\mathbf{Y})$ under the described threat model is not feasible. Instead, since our goal is to obtain a bounded leakage assurance, we leverage the existent results about the capacity of a Multi-Input Multi-Output (MIMO) channel under some specific constraints [41], to bound the mutual information, using similar techniques to [33]. Such bound is based in the fact that a MIMO channel with $K$ transmitter antennas and $c$ receiver antennas is, conceptually from an information-theoretic perspective, equivalent to a signal composed of the $K$ encoded elements of the input in PBACC, and an observation cooperatively formed by the $c$ colluding nodes, which collect and possibly process these encoded messages to infer information about the data. In this configuration, the encoded noise of the channel acts as the element that provides security, since it reduces the amount of information received by the colluding nodes. We assume for the analysis that follows that the noise is Gaussian (so the channel is an AWGN vector channel).
@@
is one of the rational interpolating terms. Without loss of generality, we assume that each entry of $X_{i}$ is a realization of a random variable with a distribution supported in the interval $D_{s}\triangleq[-s,s]$, for all $i=\{0,\dots,K-1\}$, and that the random coefficients $R_{i}\sim\mathcal{N}(0,\frac{\sigma_{n}^{2}}{T})$, where $\mathcal{N}(0,a)$ denotes the Gaussian distribution with zero mean and variance $a$. Note that, having a finite support $D_{s}$, $X_{i}$ has finite power. If we denote $\mathbf{Y}=(w(z_{i_{1}}),w(z_{i_{2}}),\dots,w(z_{i_{c}}))^{T}$, for the messages observed by the $c$ colluding adversaries $\mathcal{C}=\{i_{1},i_{2},\dots,i_{c}\}$, where $z_{i_{h}}$ are some of the Chebyshev points of second kind, we can write
$\mathbf{Y}=Q_{c}\mathbf{X}+\tilde{Q}_{c}\mathbf{R},$ (3)
where the matrices $Q_{c}$ and $\tilde{Q}_{c}$ are given by
$Q_{c}\triangleq\begin{pmatrix}q_{0}(z_{i_{1}})&\ldots&q_{K-1}(z_{i_{1}})\\ q_{0}(z_{i_{2}})&\ldots&q_{K-1}(z_{i_{2}})\\ \vdots&\ddots&\vdots\\ q_{0}(z_{i_{c}})&\ldots&q_{K-1}(z_{i_{c}})\end{pmatrix}_{c\times K},$
$\tilde{Q}_{c}\triangleq\begin{pmatrix}q_{K}(z_{i_{1}})&\ldots&q_{K+T-1}(z_{i_{1}})\\ q_{K}(z_{i_{2}})&\ldots&q_{K+T-1}(z_{i_{2}})\\ \vdots&\ddots&\vdots\\ q_{K}(z_{i_{c}})&\ldots&q_{K+T-1}(z_{i_{c}})\end{pmatrix}_{c\times T}.$
The amount of information revealed to $\mathcal{C}$ —the information leakage $I_{L}$— is defined in this work as the worst-case achievable mutual information for the colluding nodes [40]
$I_{L}\triangleq\max_{\mathcal{C}}\sup_{P_{\mathbf{X}}:||X_{i}|\leq s|,\forall i\in[K]}I(\mathbf{Y};\mathbf{X}),$
where $P_{\mathbf{X}}$ is the probability density function of $\mathbf{X}$, and the maximization applies to any $\mathcal{C}\subset[N]$. As $\|X_{i}\|\leq s$, this implies that the power $\mathbb{E}[\|X_{i}\|]^{2}\leq s^{2}$. Therefore, the latter equation can be re-written as
$I_{L}\leq\max_{\mathcal{C}}\sup_{P_{\mathbf{X}}:\mathbb{E}[\|X_{i}\|^{2}]\leq s^{2}}I(\mathbf{Y};\mathbf{X}).$ (4)
In order to bound $I_{L}$, since the noise used for privacy in PBACC is Gaussian, we consider an equivalent formulation of (3) as a MIMO channel with $K$ transmitter antennas and $c$ receiver antennas, so the input-output relation can be defined as
@@
where $H_{c\times K}$ is the channel gain matrix known by both transmitter and receiver, and $\mathbf{Z}_{c\times 1}$ is a vector of additive Gaussian noise with 0 mean. Further, let us denote the covariance matrix of the vector $\mathbf{Z}$ as $\Sigma_{\mathbf{Z}}$. Using known results on the capacity of a MIMO channel with the same power allocation constraint and correlated noise [41],
$C=\sup_{P_{\mathbf{X}}}I(\mathbf{Y};\mathbf{X})=\log_{2}|I_{c}+PH\Sigma^{-1}_{\mathbf{Z}}H^{\dagger}|,$
where $P$ is the maximum power of each transmitter antenna, $I_{c}$ is the identity matrix of order $c$ and ||$\cdot$|| ⋅ | is the determinant of a matrix. Thus, using (4) and assuming that the noise is uncorrelated, we have
$\ I_{L}\leq\max_{\mathcal{C}}\log_{2}\bigl{|}I_{c}+\frac{s^{2}T}{\sigma^{2}_{n}}\tilde{\Sigma}_{c}^{-1}\Sigma_{c}\bigr{|},$
where $\tilde{\Sigma}_{c}\triangleq\tilde{Q}_{c}\tilde{Q}_{c}^{\dagger}$ and $\Sigma_{c}\triangleq Q_{c}Q_{c}^{\dagger}$.
Since the information leakage grows with $K$, the privacy metric $\imath_{L}$ of PBACC is defined in this paper as the normalized value of $I_{L}$, namely $\imath_{L}=\frac{I_{L}}{K}$, and we will say that PBACC has $\epsilon$-privacy if $\imath_{L}<\epsilon$, where $\epsilon>0$ is the target security parameter of the scheme that represents the maximum amount of leaked information per data point that is allowed for a fixed number of colluding nodes $c$, as stated in the threat model.
To demonstrate that we meet the privacy assumptions specified in Section III-A, we will analyze the evolution of the mutual information in function of the number of colluding nodes for different values of $\sigma_{n}$. The analysis will show if our encoding provides privacy to $X$ and clarify if we can meet the privacy requirements specified in the threat model.
We start by setting up a very pessimistic scenario where we assume that if all the $N=200$ nodes of the network collude, in presence of some minimum noise $T=1$, and $s=\sigma_{n}$, they can recover all the information of $X$, being $K=10$ and $s=10^{4}$. This implies that $I_{L}=H(X)$ at most, being $H()$ the entropy of a random variable. Since the covariance matrix $\tilde{\Sigma}_{c}$ has to be inverted and it is ill-conditioned for this parameters, we regularize it using the Minimum Eigenvalue Method (very fitting for regularizing covariance matrices as per [42]). We consider that $X_{i}\sim U(-s,s)$, where $U(-a,a)$ denotes the Uniform distribution in the interval $D_{s}\triangleq[-a,a]$, to maximize the entropy. Thus, we regularize the covariance matrix until $I_{L}\approx 14.28$. At this moment, we set $T=50$ and we test the evolution of the mutual information in function of colluding nodes $c$ for $\sigma_{n}=\{10^{5},2\times 10^{5},5\times 10^{6}\}$. Fig. 2 shows the evolution of $I_{L}$, and we can clearly see that, even after starting from such a pessimistic scenario, increasing the ammount of random coefficients $T$ and the standard deviation of the noise $\sigma_{n}$ reduces the mutual information, and we can control this value to meet an $\epsilon$-security value suitable for the specific use case in study. Furthermore, we can check that the mutual information scales logarithmically, which means that new colluding nodes of the network only infer information already known by the rest of the group, which is a valuable security feature for this type of configurations.
Figure 2: Evolution of $I_{L}$ for $T=50$, $\sigma_{n}=\{10^{5},2\times 10^{5},5\times 10^{5}\}$